### v1

# Retrieval Augmented Generation for Dynamic Graph Modeling

Abstract
Dynamic graph modeling is crucial for analyzing evolving patterns in various applications. Existing approaches often integrate graph neural networks with temporal modules or redefine dynamic graph modeling as a generative sequence task. However, these methods typically rely on isolated historical contexts of the target nodes from a narrow perspective, neglecting occurrences of similar patterns or relevant cases associated with other nodes. In this work, we introduce the Retrieval-Augmented Generation for Dynamic Graph Modeling (RAG4DyG) framework, which leverages guidance from contextually and temporally analogous examples to broaden the perspective of each node. This approach presents two critical challenges: (1) How to identify and retrieve high-quality demonstrations that are contextually and temporally analogous to dynamic graph samples? (2) How can these demonstrations be effectively integrated to improve dynamic graph modeling? To address these challenges, we propose RAG4DyG, which enriches the understanding of historical contexts by retrieving and learning from contextually and temporally pertinent demonstrations. Specifically, we employ a time- and context-aware contrastive learning module to identify and retrieve relevant cases for each query sequence. Moreover, we design a graph fusion strategy to integrate the retrieved cases, thereby augmenting the inherent historical contexts for improved prediction. Extensive experiments on real-world datasets across different domains demonstrate the effectiveness of RAG4DyG for dynamic graph modeling.
1 Introduction
(a) RAG in NLP
(b) RAG in dynamic graphs
Figure 1: Illustration of RAG in NLP and dynamic graph modeling. (a) In NLP, RAG leverages pre-trained language models to encode text and retrieve semantically similar or related demonstrations, which are further concatenated to enhance the generation task. (b) Our work addresses the challenges of complex temporal and structural characteristics of dynamic graphs, incorporating RAG through time- and context-aware retrieval and graph fusion modules.
Dynamic graphs, which track the temporal evolution of nodes and edges, are essential for analyzing complex systems across domains like social networks, citation networks etc. Existing methods for dynamic graph modeling fall into two main categories: discrete-time and continuous-time approaches (Feng et al. 2024). Discrete-time methods divide the timeline into intervals, capturing the graph’s state at each point (Sankar et al. 2020), but they lack temporal granularity by overlooking event dynamics within intervals. Continuous-time approaches, on the other hand, model temporal dynamics continuously, without discretization, allowing for more detailed representations (Trivedi et al. 2019; Xu et al. 2020; Wen and Fang 2022). Our work aligns with continuous-time approaches, aiming to capture fine-grained time steps for more accurate dynamic graph modeling.
Traditional approaches for dynamic graph modeling predominantly rely on integrating graph neural networks (GNNs) with specialized temporal modules, such as recurrent neural networks (Pareja et al. 2020), self-attention (Sankar et al. 2020) and temporal point processes (Wen and Fang 2022), which update the graph representations over time. Although powerful, these methods suffer from the inherent limitations of GNNs, such as difficulty with long-term dependencies, and over-smoothing or over-squashing issues (Chen et al. 2020; Alon and Yahav 2020). Recently, a simple and effective approach called SimpleDyG (Wu, Fang, and Liao 2024) redefines dynamic graph modeling as a generative sequence modeling task and leverages the strengths of sequence models like Transformers to capture long-range dependencies within temporal sequences effectively.
Despite significant advancements in dynamic graph modeling, current approaches heavily depend on isolated historical contexts of the target nodes, limiting each node’s perspective to its ego network. The limited context makes it difficult to quickly adapt to diverse emerging patterns. For example, in a social network, when users transit to new behavioral patterns or when new users with minimal historical interactions are introduced, traditional models that focus on the historical contexts of individual nodes might not respond effectively to these changes. To overcome these limitations, we draw inspiration from the Retrieval-Augmented Generation (RAG) technique originated from the Natural Language Processing (NLP) field (Gao et al. 2023). As shown in Fig. 1(a), RAG broadens the context through the retrieval of additional demonstrations, achieving considerable success in NLP. In this work, we aim to integrate RAG into dynamic graph modeling to incorporate a broader and more relevant contextual understanding beyond the historical interactions of individual nodes. However, leveraging RAG for dynamic graph modeling presents two critical challenges.
First, how do we harvest high-quality demonstrations for dynamic graphs? This involves selecting the most contextually and temporally relevant samples to enrich the current state of a target node. Unlike RAG for NLP, where external textual data and a pre-trained language model (LM) naturally support demonstration retrieval as shown in Fig. 1(a), dynamic graph modeling requires careful consideration of the data sources and the ability to handle complex structural and temporal patterns as shown in Fig. 1(b). To address this, we introduce a novel time-aware contrastive learning strategy that leverages internal training data for demonstration retrieval. Following SimpleDyG (Wu, Fang, and Liao 2024), we regard dynamic graph modeling as a sequence modeling task and the prediction of future events as a sequence generation problem. Hence, we define a query sequence for a target node as its historical interaction sequence. Next, we automatically annotate the retrieval training pool by assessing the similarity to the query. Given the training pool, our model consists of two contrastive modules: time-aware and context-aware contrastive learning. The former utilizes a time decay function to prioritize samples that are temporally close to the query, while the latter incorporates data augmentation techniques such as masking and cropping to improve the model’s ability to retrieve complex structural patterns.
Second, how do we effectively integrate the retrieved demonstrations into the dynamic graph model? This involves fusing the retrieved demonstrations to enrich the historical context of each target node and enhance subsequent predictions. Directly concatenating the retrieved demonstrations with the query sequence not only results in a lengthy sequence unsuited for generative predictions but also overlooks the structural patterns among the retrieved demonstrations. To address this, we propose graph fusion, leveraging the graph structures inherent to the demonstrations and fusing them into a summary graph, as shown in Fig. 1(b). We then apply a GNN-based readout to learn the representation of the fused summary graph, which is prepended to the query sequence before being fed into the sequence generation model.
Our contributions can be summarized as follows. (1) We propose a novel Retrieval-Augmented Generation approach for Dynamic Graph modeling named RAG4DyG. It employs a retriever to broaden the historical interactions with contextually and temporally relevant demonstrations, leading to enhanced generative predictions on dynamic graphs. (2) We propose a time-aware contrastive learning module that incorporates temporal and structural information for demonstration retrieval. Additionally, we design a graph fusion module to effectively integrate the retrieved demonstrations, leveraging structural enhancements for subsequent predictions. (3) We conduct extensive experiments to validate our approach, demonstrating the effectiveness of RAG4DyG in various domains.
Figure 2: Overall framework of RAG4DyG. (a) Sequence modeling for dynamic graphs. (b) The retriever finds top-K temporally and contextually relevant demonstrations. (c) Graph fusion integrates the retrieved demonstrations for subsequent generation.
2 Related Work
Dynamic Graph Modeling.
Existing approaches for dynamic graphs can be categorized into discrete-time and continuous-time methods. Discrete-time methods regard a dynamic graph as a sequence of static graph snapshots captured at various time steps. Each snapshot represents the graph structure at a specific time step. These methods typically adopt GNNs to model the structural information of each snapshot, and then incorporate a sequence model (Pareja et al. 2020; Sankar et al. 2020) to capture the changes across snapshots. However, these approaches neglect fine-grained time information within a snapshot. In contrast, continuous-time methods model graph evolution as a continuous process, capturing all time steps for more precise temporal modeling. These methods often integrate GNNs with specially designed temporal modules, such as temporal random walk (Wang et al. 2021), temporal graph attention (Xu et al. 2020; Rossi et al. 2020), MLP-mixer (Cong et al. 2022) and temporal point processes (Trivedi et al. 2019; Ji et al. 2021; Wen and Fang 2022). Recently, researchers have proposed a simple and effective architecture called SimpleDyG (Wu, Fang, and Liao 2024), which reformulates the dynamic graph modeling as a sequence modeling task. Specifically, it maps the dynamic graph into a series of node sequences and feeds them into a generative sequence model. Subsequently, predicting future events can be framed as a sequence generation problem.
Retrieval Augmented Generation.
Recently, the RAG paradigm has attracted increasing attention (Gao et al. 2023). Specifically, RAG first leverages the retriever to search and extract relevant documents from some databases, which then serve as additional context to enhance the generation process. Related studies have demonstrated the great potential of RAG in various tasks such as language processing (Karpukhin et al. 2020; Jiang et al. 2023), recommendation systems (Ye et al. 2022; Dao et al. 2024), and computer vision (Liu et al. 2023; Kim et al. 2024).
In the graph modeling field, existing RAG efforts have primarily focused on static and text-attributed graphs to enhance the generation capabilities of Large Language Models, supporting graph-related tasks such as code summarization (Liu et al. 2021) and textual graph question answering (He et al. 2024; Hu et al. 2024). However, exploiting RAG techniques for dynamic graphs and graphs without textual information remains largely unexplored.
3 Preliminaries
In this section, we introduce the sequence modeling of dynamic graphs and the problem formulation.
Sequence Mapping of Dynamic Graphs.
We denote a dynamic graph as $G=(V,E,F,\mathcal{T})$ comprising a set of nodes $V$, edges $E$, a node feature matrix $F$ if available, and a time domain $\mathcal{T}$. To map a dynamic graph into sequences, we follow SimpleDyG (Wu, Fang, and Liao 2024). Specifically, let $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ denote the set of training samples, where each sample is a pair $(x_{i},y_{i})$, representing the input and output sequences for a target node $v_{i}\in V$. The input $x_{i}$ is a chronologically ordered sequence of nodes that have historically interacted with $v_{i}$, while the output $y_{i}$ is the ground truth interactions that occurs following the sequence $x_{i}$. In notations, we have
$x_{i}$ $=[\textit{hist}],v_{i},[\textit{time\_1}],v_{i}^{1,1},v_{i}^{1,2},\ldots,[\textit{time\_t}],v_{i}^{t,1},\ldots,$
[…]
Problem Formulation.
Dynamic graph modeling aims to learn a model that can predict the future interactions of a target node $v_{i}$, given its historical interactions. That is, given $x_{i}$ in Eq. (1), the task is to predict $y_{i}$ in Eq. (2).
In our RAG framework, we regard the training samples $D$ as a retrieval pool. Given a target node $v_{q}\in V$, its input sequence $x_{q}$ is referred to as the query sequence. We first retrieve $K$ demonstrations $R_{q}=\{(x_{k},y_{k})\}_{k=1}^{K}$ for each query sequence $x_{q}$ based on their contextual and temporal relevance. Next, the retrieved demonstrations $R_{q}$ are used to enrich the input sequence $x_{q}$, which encompasses the historical interactions of the target node $v_{q}$. The augmented input $\{R_{q},x_{q}\}$ is designed to enhance the predictions of future events in $y_{q}$.
4 Proposed Model: RAG4DyG
Our proposed RAG4DyG framework, depicted in Fig. 2, first trains a generative sequence model for dynamic graphs in Fig. 2(a). We use SimpleDyG (Wu, Fang, and Liao 2024) as the backbone for this task. As described in Sect. 3, a dynamic graph $G$ is represented as sequences of node interaction records, which are then used to train a Transformer-based sequence model.
Next, given a query sequence $x_{q}$ and a retrieval pool $D$, a time- and context-aware retriever is designed to retrieve demonstrations for $x_{q}$ from $D$, as shown in Fig. 2(b). The retriever fine-tunes its encoder (i.e., the sequence model) to jointly optimize two contrastive losses: a time-aware loss that employs a time decay function, and a context-aware loss that employs sequence augmentation. (Sect. 4.1)
Finally, given top-K demonstrations retrieved from the retrieval pool, we design a graph fusion module as illustrated in Fig. 2(c). Specifically, we fuse the K demonstrations into a summary graph, leveraging the graph structures inherent to the retrieved demonstrations. The summary graph is further encoded by a GNN, serving as an augmented context that is prepended to the original query sequence. The augmented sequence is subsequently input to the sequence model to predict future events. (Sect. 4.2)
4.1 Time- and Context-Aware Retriever
Unlike NLP retrievers, dynamic graph retrieval requires consideration of temporal proximity alongside contextual relevance. To address this, we propose a time- and context-aware retrieval model with two contrastive learning modules. First, we incorporate a time decay mechanism to account for temporal proximity between query and candidate sequences. Second, we use sequence augmentation to capture intrinsic contextual patterns.
Retrieval Annotation.
To facilitate contrastive training, we automatically annotate the samples in the retrieval pool $D$. For each query sequence $x_{q}$, we annotate its positive sample $x_{p}^{+}$ from the pool $D$ based on their contextual similarity. Specifically, we adopt the sequence model pre-trained in Fig. 2(a) as the encoder and apply mean pooling to obtain sequence representations. Given a query sequence $x_{q}$ and a candidate sequence $x_{p}\in D$, we define their contextual similarity as the dot product of their representations:
$s(x_{q},x_{p})=f(x_{q})^{\top}f(x_{p}),$ (3)
where $f(\cdot)$ denotes our encoder. We leave the detailed annotation process in Appendix A.
Time-aware Contrastive Learning.
Temporal information reflects the dynamic changes in historical interactions, which is crucial for dynamic graph modeling. We posit that demonstrations closer in time to the query are more relevant than those further away. Consequently, we utilize a time decay function to account for temporal proximity between the query and candidate sequences, as follows.
$\mu(t_{q},t_{p})=\exp(-\lambda|t_{q}-t_{p}|),$ (4)
where $t_{q}$ and $t_{p}$ represent the last interaction time in the query and candidate sequences [^1] In the annotated training data, the query time $t_{q}$ may precede the candidate time $t_{p}$. However, in the validation and test sets, $t_{p}$ always precedes $t_{q}$, preventing leakage from a future time., respectively. The hyper-parameter $\lambda$ controls the rate of time decay, determining how quickly the importance of interactions decreases with time. Note that $0<\mu(\cdot,\cdot)\leq 1$. By using this time decay function, we assign higher importance to the candidates that are temporally closer to the query.
To effectively capture the temporal dynamics of the graph, we incorporates temporal proximity to reweigh the contextual similarity in the contrastive loss:
$h(x_{q},x_{p})$ $=s(x_{q},x_{p})\mu(t_{q},t_{p}).$ (5)
Subsequently, we adopt in-batch negative sampling based on the following training objective:
$\mathcal{L}_{\text{tcl}}$ $=-\log\frac{\exp(h(x_{q},x_{p}^{+}))/\tau}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp(h(x_{q},x_{j}))/\tau},$ (6)
where $x_{p}^{+}$ denotes the positive sample of $x_{q}$, $N$ is the batch size, and $\tau$ is the temperature parameter.
Context-aware Contrastive Learning.
To better capture the inherent contextual pattern, we further adopt context-aware contrastive learning with data augmentations. For each sequence, we apply two types of augmentations: masking and cropping, which are widely used for sequence modeling (Devlin et al. 2019; Xie et al. 2022). The masking operator randomly replaces a portion of the tokens in the sequence with a special masking token. The cropping operator randomly deletes a contiguous subsequence from the original sequence, reducing the sequence length while preserving the temporal order of the interactions. These augmentations help the model learn robust representations and capture the inherent structural information of the sequence by focusing on its different parts.
We treat two augmented views of the same sequence as positive pairs, and those of different sequences as negative pairs. Given a sequence $x_{q}$ and its two distinct augmented views $x^{\prime}_{q}$ and $x^{\prime\prime}_{q}$, the contrastive loss is defined as:
$\mathcal{L}_{\text{ccl}}=-\log\frac{\exp(s(x^{\prime}_{q},x^{\prime\prime}_{q})/\tau)}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp s(x^{\prime}_{q},x^{\prime}_{j})/\tau},$ (7)
where $\tau$ is the temperature, $N$ is the batch size and $\mathds{1}$ is an indicator function.
Training and Inference for Retrieval.
The training objective of our retrieval model is defined as:
$\mathcal{L}_{\text{ret}}=\mathcal{L}_{\text{tcl}}+\alpha\mathcal{L}_{\text{ccl}}$ (8)
where $\alpha$ is a coefficient that controls the balance between the two losses.
During testing, we utilize the updated sequence model to extract sequence representations and perform demonstration ranking based on the contextual similarity between the query and candidates.
4.2 Graph Fusion-based Generator
After retrieval, we obtain the top-K demonstrations $R_{q}=\{(x_{k},y_{k})\}_{k=1}^{K}$ for the query $x_{q}$. A straightforward approach is concatenating them with the query sequence. However, this can lead to a lengthy context that limits the model’s prediction capabilities. More importantly, it neglects the structural patterns among these demonstrations. Thus, we first fuse the demonstrations into a summary graph, process it using a GNN, and then prepend the graph readout from the GNN to the query for subsequent generation.
Graph Fusion.
[…]
$e_{\text{fus}}$ $=\mathtt{MeanPooling}(\text{GCN}(G_{\text{fus}})),$ (9)
$\tilde{x}_{q}$ $=[e_{\text{fus}}\parallel x_{q}],$ (10)
where $e_{\text{fus}}$ is the fused graph representation, and $\tilde{x}_{q}$ is the retrieval-augmented sequence. The augmented sequence is fed into the sequence model, which generates future interactions.
Training and inference.
We adopt the same sequence model with the same training objective (Wu, Fang, and Liao 2024) as in Fig. 2(a). During training, we freeze the parameters of the sequence model, except for the output layer which is updated along with the GCN parameters used for graph fusion.
During testing, we first apply the retriever model to retrieve top-K demonstrations for each query as introduced in Sect. 4.1. Then we perform graph fusion on these demonstrations and concatenate the fused graph representation with the query sequence as illustrated in Eqs. (9) and (10). The concatenated sequence is subsequently fed into the trained sequence model for link prediction.
5 Experiment
In this section, we empirically evaluate the proposed model RAG4DyG compared to state-of-the-art methods and conduct a detailed analysis of the performance.
5.1 Experimental Setup
Datasets.
We evaluate the performance of the proposed model on three datasets from different domains including the communication network UCI (Panzarasa, Opsahl, and Carley 2009), the citation network Hepth (Leskovec, Kleinberg, and Faloutsos 2005), and the multi-turn task-oriented conversation dataset MMConv (Liao et al. 2021). We follow the same dataset preprocessing as SimpleDyG (Wu, Fang, and Liao 2024). It is important to note that the ML-10M dataset was not used because it is not suitable for RAG; we observed no performance enhancement even when the query was augmented with ground-truth demonstrations. We summarize the statistics of the three datasets in Table 1.
Dataset UCI Hepth MMConv
Domain Social Citation Conversation
# Nodes 1,781 4,737 7,415
# Edges 16,743 14,831 91,986
Table 1: Dataset statistics.
Method UCI Hepth MMConv
Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard
DySAT 0.009±0.003 0.010±0.003 0.010±0.001 0.008±0.004 0.007±0.002 0.005±0.001 0.108±0.089 0.102±0.085 0.095±0.080
EvolveGCN 0.072±0.046 0.064±0.045 0.032±0.026 0.008±0.002 0.009±0.004 0.007±0.002 0.050±0.015 0.051±0.021 0.032±0.017
DyRep 0.009±0.008 0.011±0.018 0.010±0.005 0.009±0.006 0.031±0.024 0.010±0.006 0.156±0.054 0.140±0.057 0.067±0.025
JODIE 0.018±0.019 0.022±0.023 0.012±0.009 0.010±0.008 0.031±0.021 0.011±0.008 0.052±0.039 0.041±0.016 0.032±0.022
TGAT 0.022±0.004 0.061±0.007 0.020±0.002 0.011±0.007 0.034±0.023 0.011±0.006 0.118±0.004 0.089±0.033 0.058±0.021
TGN 0.014±0.010 0.041±0.017 0.011±0.003 0.011±0.006 0.030±0.012 0.008±0.001 0.085±0.050 0.096±0.068 0.066±0.038
TREND 0.083±0.015 0.067±0.010 0.039±0.020 0.010±0.008 0.031±0.003 0.010±0.002 0.134±0.030 0.116±0.020 0.060±0.018
GraphMixer 0.097±0.019 0.104± 0.013 0.042 ±0.005 0.009±0.002 0.011±0.008 0.010±0.003 0.206±0.001 0.172±0.029 0.085±0.016
SimpleDyG 0.109±0.014 0.104±0.010 0.092±0.014 0.013±0.006 0.035±0.014 0.013±0.006 0.170±0.010 0.184±0.012 0.169±0.010
RAG4DyG 0.111±0.013 0.122±0.014 0.097±0.010 0.019±0.002 0.045±0.003 0.019±0.002 0.194±0.005 0.208±0.005 0.194±0.005
Table 2: Performance comparison for dynamic link prediction. (Best results are bolded; runners-up are underlined.)
Baselines.
We compare RAG4DyG with the state-of-the-art approaches of dynamic graph modeling. (1) discrete-time approaches: DySAT (Sankar et al. 2020) and EvolveGCN (Pareja et al. 2020); (2) continuous-time approaches: DyRep (Trivedi et al. 2019), JODIE (Kumar, Zhang, and Leskovec 2019), TGAT (Xu et al. 2020), TGN (Rossi et al. 2020), TREND (Wen and Fang 2022), GraphMixer (Cong et al. 2022) and SimpleDyG (Wu, Fang, and Liao 2024). The details of the baselines can be found in Appendix B.
Implementation Details.
Following the method outlined in (Cong et al. 2022; Wu, Fang, and Liao 2024), we represent the dynamic graph as an undirected graph. We split all datasets into training, validation, and test sets based on temporal sequence same as SimpleDyG (Wu, Fang, and Liao 2024). Given ${T}$ timesteps in each dataset, the data at the final timestep ${T}$ is used as the testing set, the data at ${T-1}$ is served as the validation set, and the remaining data from earlier timesteps is used for training. All training data including the retrieval pool for the retriever and generator is drawn from this training data split. For retrieval augmented generation model training, we first train SimpleDyG without augmentation using the reported parameter in this paper. Then we fix the parameters of SimipleDyG except for the last linear layer and fine-tune them with the GCN model. The number of GCN layers in the generator model is 1 for all datasets. We repeat each experiment 10 times and report the average results along with the standard deviation. The number of demonstrations is 7 for all datasets. Hyper-parameter tuning is performed on the validation set and we provide the settings in Appendix D.
Evaluation Metrics.
We evaluate the performance upon three metrics Recall@5, NDCG@5 and Jaccard (Wu, Fang, and Liao 2024). Recall@5 and NDCG@5 are widely used in ranking tasks measuring the quality of top-ranked predictions (Wang et al. 2019). Recall@5 assesses the proportion of relevant nodes in the top 5 predictions, while NDCG@5 evaluates the ranking quality by considering the positions of the relevant nodes. Jaccard (Jaccard 1901) measures the similarity between the predicted and ground truth sequences by comparing the intersection and union of the two sets.
5.2 Performance Comparison
We evaluate the performance of RAG4DyG for the dynamic link prediction task, and the results compared to the state-of-the-art baselines are presented in Table 2. We make the following observations.
The proposed RAG4DyG generally outperforms all baselines across different datasets under the three metrics. In particular, compared to SimpleDyG, which is also our backbone, RAG4DyG consistently shows superior performance, highlighting the effectiveness of our retrieval-augmented generation framework. Note that GraphMixer performs slightly better in Recall@5 on the MMConv dataset, but its significantly lower performance in NDCG@5 and Jaccard indicates that its predictions are not ranked optimally or maintaining the overall set integrity compared to RAG4DyG. This indicates that RAG4DyG can better model the temporal and contextual relationships due to the specific design of the retriever and generator.
Figure 3: Ablation study for retrieval results.
Second, RAG4DyG exhibits significant advantages in inductive scenarios such as the Hepth dataset. This setting is particularly challenging because it involves nodes not appearing during the training phase, requiring the model to generalize to entirely new structures and relationships. RAG4DyG’s success is attributed to its retrieval-augmented mechanism, which enhances the model’s ability to generalize by providing rich contextual information relevant to the new, unseen nodes. Unlike models that rely solely on the immediate neighborhood or predefined structures, RAG4DyG dynamically adapts to the new nodes, ensuring that the predictions are guided by the most relevant and similar historical data.
5.3 Model Analysis
We analyze the behavior of our model RAG4DyG in several aspects, including an ablation study, an investigation of the effectiveness of different retrieval methods, and an analysis of parameter sensitivity.
Ablation Study.
To evaluate the effectiveness of different modules in the retrieval model, we compare RAG4DyG with two variants w/o CCL and w/o Decay which exclude the context-aware contrastive learning and time decay component in the retrieval model. We evaluate the performance for both retrieval and link prediction tasks. We use HR@k (Hit Ratio@k) metrics for the retrieval model, measuring the proportion of cases where at least one of the top-k retrieved items is relevant. As shown in Fig. 3 and 4, the full model outperforms the two variants, underscoring the benefits of incorporating context-aware contrastive learning and time decay modulation.
Figure 4: Ablation study for link prediction results.
Effect of Different Retrieval Methods.
To further investigate the effectiveness of the retrieval model, we compare our model with two different retrieval methods, namely, BM25 and Jaccard, in Table 3 and 4. In Table 4, we only report NDCG@5 and present the remaining metrics in Appendix C. BM25 (Fang, Tao, and Zhai 2004) is an extension of the Term Frequency-Inverse Document Frequency (TF-IDF) model, which calculates a relevance score between the query sequence and each candidate sequence in the retrieval pool. The relevance score is derived from the occurrence frequency of the nodes in the query and the retrieval pool. Jaccard (Jaccard 1901) measures the similarity between two sets by comparing the size of their intersection to the size of their union. Note that in the citation dataset Hepth, all queries in the test set only contain unseen target nodes that never appear in the retrieval pool and have no historical interactions. As a result, the BM25 and Jaccard scores between the queries and the candidates in the retrieval pool are always zeros. On the other hand, our retrieval model is trained based on the sequence representations. For a query sequence containing only the target node, we can still obtain its representation using the sequence model trained for the retrieval model, and further calculate its contextual similarity with the candidate sequences in the retrieval pool.
Method UCI Hepth MMConv
HR@1 HR@3 HR@1 HR@3 HR@1 HR@3
BM25 0.100 0.136 - - 0.025 0.233
Jaccard 0.100 0.109 - - 0.007 0.213
RAG4DyG 0.096 0.153 0.025 0.049 0.034 0.168
“-” denotes that the method is unable to perform retrieval. The reason is explained in the corresponding description of this table in Sect. 5.3.
Table 3: Retrieval performance of various retrieval methods.
Method UCI Hepth MMConv
BM25 0.121±0.009 - 0.207±0.003
Jaccard 0.$1$ 0.207±0.005
RAG4DyG 0.122±0.014 0.045±0.003 0.208±0.005
GroundTruth 0.129±0.010 0.062±0.007 0.224±0.008
See the note in Table 3 for the explanation of “-”.
Table 4: Generative accuracy of various retrieval methods.
In Table 3, we analyze the retrieval performance of different retrieval methods. Our retrieval model shows better performance than other retrieval strategies. Notably, in inductive scenarios like the Hepth dataset, BM25 and Jaccard fail to work with new query nodes lacking historical interactions. In contrast, our model can handle them effectively and achieve solid performance. The high $HR@3$ performance of BM25 and Jaccard on the MMConv dataset can be attributed to the nature of the dialogue dataset, where temporal order is less critical, and certain nodes associated with specific slot values are more discriminative.
Table 4 shows the generative performance of different retrieval methods in the dynamic link prediction task. During testing, we apply the retrieval results obtained from different retrieval methods. We also train a model using the ground-truth retrieval results for a more comprehensive comparison. The “GroundTruth” row represents an upper bound on the performance when using ground-truth retrieval results on the testing data, which, as expected, provides the highest performance metrics. Generally speaking, all retrieval methods show better performance compared to the backbone SimpleDyG without using RAG, demonstrating the effectiveness of the RAG technique for dynamic graph modeling. Our method performs better compared to other retrieval strategies, indicating the effectiveness of contrastive learning in the retrieval model.
Effect of the Number of Demonstrations K.
To investigate the influence of the number of demonstrations, we conduct experiments across varying values $K\in\{1,3,5,7,9\}$. As shown in Fig. 5, a higher number of $K$ yields better prediction performance, that’s because more demonstrations provide richer contextual information, especially in the UCI dataset. However, including too many cases may introduce more noise, which can harm the performance.
Figure 5: Effect of the number of demonstrations K.
Effect of Different Fusion Strategies.
To further investigate the effectiveness of the fusion strategy for the top-K demonstrations, we conduct experiments with different fusion strategies in Table 5 for K $=7$. We report NDCG@5 in Table 4 and present the other metrics in Appendix C. “Concatenation” denotes we directly concatenate the sequences of retrieved demonstrations and prepend them with the query sample sequence and then feed it into the pre-trained SimpleDyG model. “MLP” means we do not consider the graph structure of the demonstrations and replace the graph fusion as an MLP layer (we set the number of the MLP layer as 2). By using the MLP layer, We map the concatenated demonstrations into shorter m-dimensional embeddings (we empirically set m to be 15) and then concatenate it with the query sample. Like graph fusion, we only fine-tune the parameters of the MLP and output layer. The results in Table 5 show that directly concatenating the retrieved demonstrations with the query sample leads to lower performance compared with other strategies. This is because simple concatenation introduces a lengthy context, which can overwhelm the model with irrelevant information, and it neglects the structural relationships inherent in the demonstrations. The “MLP” strategy improves upon this by mapping the concatenated demonstrations into a shorter feature space, effectively reducing noise and emphasizing more relevant features. This approach yields better results than simple concatenation but still falls short compared to the “GraphFusion” strategy. The superior performance of the “GraphFusion” strategy highlights the importance of considering both the content and the structure of the demonstrations in the fusion process.
Fusion strategy UCI Hepth MMConv
Concatenation 0.036±0.018 0.007±0.002 0.110±0.030
MLP 0.106±0.017 0.015±0.002 0.167±0.006
GraphFusion 0.122±0.014 0.045±0.003 0.208±0.005
Table 5: Effect of different fusion strategies.
6 Conclusion
In this work, we introduced a novel retrieval augmented framework RAG4DyG for dynamic graph modeling, addressing the limitations of existing approaches that often rely on narrow historical contexts. We leverage the RAG technique to broaden the context by incorporating relevant auxiliary information by harvesting high-quality demonstrations for dynamic graph samples and effectively utilizing this demonstration to enhance dynamic graph modeling. Our proposed solution involves a time-aware contrastive learning model to identify and retrieve temporally pertinent cases for each query sequence, coupled with a graph fusion strategy to integrate the inherent historical context with extended temporal contexts. Extensive experiments on real-world datasets across different domains demonstrated the superior performance of RAG4DyG in dynamic graph modeling.
Appendix
Appendix A Retrieval Data Annotation
To facilitate retrieval model training, we regard the samples in the training dataset as our retrieval pool $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ where each pair $(x_{i},y_{i})$ represents the historical sequence and its corresponding target sequence. Specifically, $x_{i}$ is the input sequence before the last time step and $y_{i}$ is the output sequence at the last time step. We annotate demonstrations based on the Jaccard similarity of $y_{i}$ among all the pairs in $D$.
$r(y_{i},y_{j})=\frac{|y_{i}\cap y_{j}|}{|y_{i}\cup y_{j}|}$ (11)
To control the quality of annotated data, we set a threshold of 0.8 to select highly similar demonstrations for each sample. These filtered annotations are then used to train the retriever model. The number of training samples for UCI, Hepth and MMConv datasets are 9,578, 8,250 and 10,762, respectively.
Appendix B More Details about Baselines
The details of the baselines are as follows:
  • DySAT (Sankar et al. 2020) utilizes self-attention mechanisms to capture both structural and temporal patterns in dynamic graphs through discrete-time snapshots.
  • EvolveGCN (Pareja et al. 2020) leverages recurrent neural networks to evolve the graph convolutional network parameters over discrete time steps.
  • DyRep (Trivedi et al. 2019) models dynamic graphs in continuous time by incorporating both temporal point processes and structural dynamics to capture interactions and node dynamics.
  • JODIE (Kumar, Zhang, and Leskovec 2019) focuses on user and item embedding trajectories over continuous time, predicting future interactions by modeling user and item embeddings jointly.
  • TGAT (Xu et al. 2020) employs temporal graph attention layers and time encoding to capture temporal dependencies and structural information for dynamic graphs.
  • TGN (Rossi et al. 2020) combines GNNs with memory modules to maintain node states over continuous time, effectively learning from dynamic interactions.
  • TREND (Wen and Fang 2022) integrates temporal dependencies based on Hawkes process and GNN to learn the dynamics of graphs.
  • GraphMixer (Cong et al. 2022) introduces a novel architecture that leverages MLP-mixer to learn link-encoder and node encoder for evolving graphs in continuous time.
  • SimpleDyG (Wu, Fang, and Liao 2024) reformulated the dynamic graph modeling as a sequence modeling task and mapped the dynamic interactions of target nodes as sequences with specially designed tokens. It simplifies dynamic graph modeling without complex architectural changes to effectively capture temporal dynamics.
Appendix C More Experimental Results
For the effect of different retrieval methods for generative performance in Sec. 5.3, we report the NDCG@5 score in Table 4. Here we report the remaining metrics in Table 6. The results show that RAG4DyG performs better than BM25 and Jaccard on Recall@5 and Jaccard metrics, which are consistent with the performance of NDCG@5.
Method UCI Hepth MMConv
Recall@5 Jaccard Recall@5 Jaccard Recall@5 Jaccard
BM25 0.111±0.007 0.093±0.004 - - 0.193±0.003 0.192±0.003
Jaccard 0.104±0.009 0.088±0.010 - - 0.193±0.004 0.192±0.004
RAG4DyG 0.111±0.013 0.097±0.010 0.019±0.002 0.019±0.002 0.194±0.005 0.194 ±0.005
GroundTruth 0.121±0.010 0.107±0.012 0.028±0.004 0.028±0.004 0.210±0.008 0.208±0.008
Table 6: Generative performance of retrieval methods.
For the effect of different fusion strategies in Sec. 5.3, the performance of the remaining metrics is shown in Table 7. Similarly, the “GraphFusion” strategy outperforms direct concatenation and “MLP”, indicating the effectiveness of the graph structure among demonstrations.
Fusion strategy UCI Hepth MMConv
Recall@5 Jaccard Recall@5 Jaccard Recall@5 Jaccard
Concatenation 0.033±0.019 0.029±0.016 0.001±0.002 0.002±0.002 0.103±0.031 0.103±0.031
MLP 0.102±0.018 0.089±0.016 0.006±0.002 0.006±0.002 0.156±0.006 0.153±0.006
GraphFusion 0.111±0.013 0.097±0.010 0.019±0.002 0.019±0.002 0.194±0.005 0.194 ±0.005
Table 7: Effect of different fusion strategies.
Appendix D Hyper-parameter Settings
We run all the experiments on an NVIDIA GPU L40 with the same data splitting and follow the hyper-parameter settings reported in SimpleDyG (Wu, Fang, and Liao 2024) for all the baselines. For our RAG4DyG method, we report the hyper-parameters as follows.
The time decay rate $\lambda$ for the retrieval model in Eq. (4) was fine-tuned according to the time granularity of different datasets, with days for the UCI dataset, months for the Hepth dataset, and turns for the MMConv dataset. We explored a range of values $\lambda=\{1e-4,1e-3,1e-2,1e-1,1,10,100\}$, ultimately selecting $\lambda=1e-4$ for UCI, $\lambda=0.1$ for Hepth, and $\lambda=10$ for MMConv. The coefficient $\alpha$ in the loss function (Eq. (8) was tuned across $\{0.2,0.4,0.6,0.8,1\}$, resulting in final values of $\alpha=1$ for UCI and MMConv dataset, $\alpha=0.4$ for Hepth. Additional parameter settings for the three datasets are as follows: the temperature $\tau$ in the two contrastive learning tasks is set to $\tau=\{0.1,0.1,0.1\}$; the batch size of the retriever model is $N=\{64,128,128\}$; and the masking and cropping portions in context-aware contrastive learning are $\{0.8,0.8,0.8\}$ and $\{0.4,0.6,0.6\}$, respectively.

### v2

@@
 Retrieval Augmented Generation for Dynamic Graph Modeling
-Abstract
-Dynamic graph modeling is crucial for analyzing evolving patterns in various applications.
-Existing approaches often integrate graph neural networks with temporal modules or redefine dynamic graph modeling as a generative sequence task.
-However, these methods typically rely on isolated historical contexts of the target nodes from a narrow perspective, neglecting occurrences of similar patterns or relevant cases associated with other nodes.
-In this work, we introduce the Retrieval-Augmented Generation for Dynamic Graph Modeling (RAG4DyG) framework, which leverages guidance from contextually and temporally analogous examples to broaden the perspective of each node.
-This approach presents two critical challenges: (1) How to identify and retrieve high-quality demonstrations that are contextually and temporally analogous to dynamic graph samples?
-(2) How can these demonstrations be effectively integrated to improve dynamic graph modeling?
-To address these challenges, we propose RAG4DyG, which enriches the understanding of historical contexts by retrieving and learning from contextually and temporally pertinent demonstrations.
-Specifically, we employ a time- and context-aware contrastive learning module to identify and retrieve relevant cases for each query sequence.
-Moreover, we design a graph fusion strategy to integrate the retrieved cases, thereby augmenting the inherent historical contexts for improved prediction.
-Extensive experiments on real-world datasets across different domains demonstrate the effectiveness of RAG4DyG for dynamic graph modeling.
-1 Introduction
+Abstract.
+Modeling dynamic graphs, such as those found in social networks, recommendation systems, and e-commerce platforms, is crucial for capturing evolving relationships and delivering relevant insights over time.
+Traditional approaches primarily rely on graph neural networks with temporal components or sequence generation models, which often focus narrowly on the historical context of target nodes.
+This limitation restricts the ability to adapt to new and emerging patterns in dynamic graphs.
+To address this challenge, we propose a novel framework, Retrieval-Augmented Generation for Dynamic Graph modeling (RAG4DyG), which enhances dynamic graph predictions by incorporating contextually and temporally relevant examples from broader graph structures.
+Our approach includes a time- and context-aware contrastive learning module to identify high-quality demonstrations and a graph fusion strategy to effectively integrate these examples with historical contexts.
+The proposed framework is designed to be effective in both transductive and inductive scenarios, ensuring adaptability to previously unseen nodes and evolving graph structures.
+Extensive experiments across multiple real-world datasets demonstrate the effectiveness of RAG4DyG in improving predictive accuracy and adaptability for dynamic graph modeling.
+The code and datasets are publicly available at https://github.com/YuxiaWu/RAG4DyG.
+Retrieval-augmented generation, graph neural networks, dynamic graph modeling
+journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18,2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729958isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Information retrievalccs: Computing methodologies Artificial intelligence
+1. Introduction
 (a) RAG in NLP
 (b) RAG in dynamic graphs
-Figure 1: Illustration of RAG in NLP and dynamic graph modeling.
+Figure 1.
+Illustration of RAG in NLP and dynamic graph modeling.
 (a) In NLP, RAG leverages pre-trained language models to encode text and retrieve semantically similar or related demonstrations, which are further concatenated to enhance the generation task.
 (b) Our work addresses the challenges of complex temporal and structural characteristics of dynamic graphs, incorporating RAG through time- and context-aware retrieval and graph fusion modules.
-Dynamic graphs, which track the temporal evolution of nodes and edges, are essential for analyzing complex systems across domains like social networks, citation networks etc.
-Existing methods for dynamic graph modeling fall into two main categories: discrete-time and continuous-time approaches (Feng et al.
-2024).
-Discrete-time methods divide the timeline into intervals, capturing the graph’s state at each point (Sankar et al.
-2020), but they lack temporal granularity by overlooking event dynamics within intervals.
-Continuous-time approaches, on the other hand, model temporal dynamics continuously, without discretization, allowing for more detailed representations (Trivedi et al.
-2019; Xu et al.
-2020; Wen and Fang 2022).
-Our work aligns with continuous-time approaches, aiming to capture fine-grained time steps for more accurate dynamic graph modeling.
-Traditional approaches for dynamic graph modeling predominantly rely on integrating graph neural networks (GNNs) with specialized temporal modules, such as recurrent neural networks (Pareja et al.
-2020), self-attention (Sankar et al.
-2020) and temporal point processes (Wen and Fang 2022), which update the graph representations over time.
-Although powerful, these methods suffer from the inherent limitations of GNNs, such as difficulty with long-term dependencies, and over-smoothing or over-squashing issues (Chen et al.
-2020; Alon and Yahav 2020).
-Recently, a simple and effective approach called SimpleDyG (Wu, Fang, and Liao 2024) redefines dynamic graph modeling as a generative sequence modeling task and leverages the strengths of sequence models like Transformers to capture long-range dependencies within temporal sequences effectively.
-Despite significant advancements in dynamic graph modeling, current approaches heavily depend on isolated historical contexts of the target nodes, limiting each node’s perspective to its ego network.
-The limited context makes it difficult to quickly adapt to diverse emerging patterns.
-For example, in a social network, when users transit to new behavioral patterns or when new users with minimal historical interactions are introduced, traditional models that focus on the historical contexts of individual nodes might not respond effectively to these changes.
-To overcome these limitations, we draw inspiration from the Retrieval-Augmented Generation (RAG) technique originated from the Natural Language Processing (NLP) field (Gao et al.
-2023).
-As shown in Fig.
-1(a), RAG broadens the context through the retrieval of additional demonstrations, achieving considerable success in NLP.
-In this work, we aim to integrate RAG into dynamic graph modeling to incorporate a broader and more relevant contextual understanding beyond the historical interactions of individual nodes.
-However, leveraging RAG for dynamic graph modeling presents two critical challenges.
-First, how do we harvest high-quality demonstrations for dynamic graphs?
-This involves selecting the most contextually and temporally relevant samples to enrich the current state of a target node.
-Unlike RAG for NLP, where external textual data and a pre-trained language model (LM) naturally support demonstration retrieval as shown in Fig.
-1(a), dynamic graph modeling requires careful consideration of the data sources and the ability to handle complex structural and temporal patterns as shown in Fig.
-1(b).
-To address this, we introduce a novel time-aware contrastive learning strategy that leverages internal training data for demonstration retrieval.
-Following SimpleDyG (Wu, Fang, and Liao 2024), we regard dynamic graph modeling as a sequence modeling task and the prediction of future events as a sequence generation problem.
-Hence, we define a query sequence for a target node as its historical interaction sequence.
-Next, we automatically annotate the retrieval training pool by assessing the similarity to the query.
-Given the training pool, our model consists of two contrastive modules: time-aware and context-aware contrastive learning.
-The former utilizes a time decay function to prioritize samples that are temporally close to the query, while the latter incorporates data augmentation techniques such as masking and cropping to improve the model’s ability to retrieve complex structural patterns.
-Second, how do we effectively integrate the retrieved demonstrations into the dynamic graph model?
-This involves fusing the retrieved demonstrations to enrich the historical context of each target node and enhance subsequent predictions.
-Directly concatenating the retrieved demonstrations with the query sequence not only results in a lengthy sequence unsuited for generative predictions but also overlooks the structural patterns among the retrieved demonstrations.
-To address this, we propose graph fusion, leveraging the graph structures inherent to the demonstrations and fusing them into a summary graph, as shown in Fig.
-1(b).
-We then apply a GNN-based readout to learn the representation of the fused summary graph, which is prepended to the query sequence before being fed into the sequence generation model.
-Our contributions can be summarized as follows.
-(1) We propose a novel Retrieval-Augmented Generation approach for Dynamic Graph modeling named RAG4DyG.
-It employs a retriever to broaden the historical interactions with contextually and temporally relevant demonstrations, leading to enhanced generative predictions on dynamic graphs.
-(2) We propose a time-aware contrastive learning module that incorporates temporal and structural information for demonstration retrieval.
-Additionally, we design a graph fusion module to effectively integrate the retrieved demonstrations, leveraging structural enhancements for subsequent predictions.
-(3) We conduct extensive experiments to validate our approach, demonstrating the effectiveness of RAG4DyG in various domains.
-Figure 2: Overall framework of RAG4DyG.
+Dynamic graph modeling is essential for understanding and predicting evolving interactions across various applications such as social networks (Deng et al., 2019; Sun et al., 2022), personalized recommendations(Song et al., 2019; Wu et al., 2022; Tang et al., 2023), and Web-based services (Han et al., 2021; Duan et al., 2024).
+These applications require effective handling of the temporal and contextual dynamics inherent in evolving graph structures to provide accurate and timely insights.
+For example, in a recommendation system (Tang et al., 2023), capturing user behavioral shifts over time can enhance the personalization of content, while in social networks (Sun et al., 2022), understanding interaction patterns can improve engagement and fraud detection mechanisms.
+Existing dynamic graph modeling methods generally fall into discrete-time and continuous-time approaches (Feng et al., 2024).
+Discrete-time models capture graph snapshots at specific intervals, providing a simplified representation of the graph’s evolution but often failing to account for fine-grained temporal dynamics (Sankar et al., 2020; Pareja et al., 2020).
+Continuous-time models, such as DyRep (Trivedi et al., 2019), TGAT (Xu et al., 2020), and TREND (Wen and Fang, 2022), model events as they occur, offering a more granular perspective and better capturing event-driven dynamics.
+These models commonly rely on graph neural networks (GNNs) integrated with temporal mechanisms such as recurrent neural networks (Pareja et al., 2020), self-attention (Sankar et al., 2020), and temporal point processes (Wen and Fang, 2022) to update graph representations over time.
+Despite their strengths, GNN-based methods struggle with long-term dependencies and issues such as over-smoothing and over-squashing (Chen et al., 2020; Alon and Yahav, 2020).
+Recently, SimpleDyG (Wu et al., 2024) redefined dynamic graph modeling as a generative sequence modeling task, leveraging Transformers to effectively capture long-range dependencies within temporal sequences.
+However, the reliance on localized historical interactions still limits the ability to generalize across different contexts and adapt to emerging patterns.
+To address these limitations, the Retrieval-Augmented Generation (RAG) framework from the Natural Language Processing (NLP) domain (Gao et al., 2023) offers a promising approach.
+RAG has the potential to broaden the contextual understanding of dynamic graphs by retrieving and incorporating relevant examples from across the graph’s temporal and contextual space, as illustrated in Figure 1.
+However, adopting RAG for dynamic graph modeling presents two major challenges: (1) Selecting high-quality demonstrations, and (2) effectively integrating the retrieved demonstrations.
+Identifying contextually and temporally relevant demonstrations is challenging because existing retrieval methods, such as BM25 and other matching-based schemes (Robertson and Jones, 1976) primarily rely on historical interaction similarities and struggle with inductive scenarios where nodes lack historical interactions.
+Moreover, effective integration of retrieved demonstrations is another challenge, as simply concatenating them with the query sequence can lead to overly lengthy inputs and overlook underlying structural patterns.
+To overcome these challenges, we propose a novel Retrieval-Augmented Generation for Dynamic Graph modeling (RAG4DyG) framework.
+It integrates a time- and context-aware contrastive learning strategy that evaluates historical interaction sequences to identify relevant examples and a graph fusion module to construct a summary graph from the retrieved samples.
+The contrastive learning strategy incorporates a time decay function to prioritize temporally relevant samples, while context-aware augmentation techniques such as masking and cropping enhance the model’s ability to capture complex structural patterns.
+The graph fusion module applies a GNN-based readout mechanism to enrich the representation before feeding it into the sequence generation model.
+These solutions empower RAG4DyG to effectively leverage retrieved demonstrations to enhance dynamic graph modeling.
+Through extensive experimentation on various real-world datasets, we demonstrate that RAG4DyG outperforms state-of-the-art methods in both transductive and inductive scenarios, offering improved accuracy and adaptability in dynamic graph scenarios.
+In transductive settings, where test nodes have appeared during training, our model effectively leverages historical data to refine predictions.
+In inductive settings, involving previously unseen nodes, the retrieval mechanism enables the model to generalize by providing relevant contextual examples as guidance.
+To sum up, our main contributions are as follows.
+(1) We propose a novel retrieval-augmented generation approach for dynamic graph modeling named RAG4DyG, which employs a retriever to broaden historical interactions with contextually and temporally relevant demonstrations.
+(2) We introduce a time- and context-aware contrastive learning module that incorporates temporal and structural information for demonstration retrieval and a graph fusion module to effectively integrate retrieved demonstrations.
+(3) We conduct extensive experiments to validate our approach, demonstrating the effectiveness of RAG4DyG across various domains.
+Figure 2.
+Overall framework of RAG4DyG.
 (a) Sequence modeling for dynamic graphs.
 (b) The retriever finds top-K temporally and contextually relevant demonstrations.
-(c) Graph fusion integrates the retrieved demonstrations for subsequent generation.
-2 Related Work
+(c) Graph fusion integrates the retrieved demonstrations for the subsequent generation.
+2. Related Work
 Dynamic Graph Modeling.
@@
 Each snapshot represents the graph structure at a specific time step.
-These methods typically adopt GNNs to model the structural information of each snapshot, and then incorporate a sequence model (Pareja et al.
-2020; Sankar et al.
-2020) to capture the changes across snapshots.
+These methods typically adopt GNNs to model the structural information of each snapshot, and then incorporate a sequence model (Pareja et al., 2020; Sankar et al., 2020) to capture the changes across snapshots.
 However, these approaches neglect fine-grained time information within a snapshot.
 In contrast, continuous-time methods model graph evolution as a continuous process, capturing all time steps for more precise temporal modeling.
-These methods often integrate GNNs with specially designed temporal modules, such as temporal random walk (Wang et al.
-2021), temporal graph attention (Xu et al.
-2020; Rossi et al.
-2020), MLP-mixer (Cong et al.
-2022) and temporal point processes (Trivedi et al.
-2019; Ji et al.
-2021; Wen and Fang 2022).
-Recently, researchers have proposed a simple and effective architecture called SimpleDyG (Wu, Fang, and Liao 2024), which reformulates the dynamic graph modeling as a sequence modeling task.
+These methods often integrate GNNs with specially designed temporal modules, such as temporal random walk (Wang et al., 2021), temporal graph attention (Xu et al., 2020; Rossi et al., 2020), MLP-mixer (Cong et al., 2022) and temporal point processes (Trivedi et al., 2019; Ji et al., 2021; Wen and Fang, 2022).
+Recently, researchers have proposed a simple and effective architecture called SimpleDyG (Wu et al., 2024), which reformulates dynamic graph modeling as a sequence modeling task.
 Specifically, it maps the dynamic graph into a series of node sequences and feeds them into a generative sequence model.
 Subsequently, predicting future events can be framed as a sequence generation problem.
+However, while these methods provide valuable insights, they are often limited in their ability to adapt to new and evolving patterns in dynamic graphs.
+Our work distinguishes itself from prior dynamic graph learning methods through two key innovations.
+First, while existing approaches predominantly focus on localized temporal contexts or the historical interactions of target nodes, our proposed RAG4DyG employs retrieval-augmented generation mechanisms to retrieve and integrate broader contextual signals from the entire dynamic graph.
+This approach facilitates a more comprehensive understanding of dynamic interactions, uncovering complex patterns beyond the immediate historical scope of individual nodes.
+Second, RAG4DyG incorporates a time- and context-aware contrastive learning module for retrieving similar demonstrations, along with a graph fusion strategy to integrate them with the query sequence, enhancing adaptability to new patterns and evolving graph structures.
 Retrieval Augmented Generation.
-Recently, the RAG paradigm has attracted increasing attention (Gao et al.
-2023).
+Recently, the RAG paradigm has attracted increasing attention (Gao et al., 2023).
 Specifically, RAG first leverages the retriever to search and extract relevant documents from some databases, which then serve as additional context to enhance the generation process.
-Related studies have demonstrated the great potential of RAG in various tasks such as language processing (Karpukhin et al.
-2020; Jiang et al.
-2023), recommendation systems (Ye et al.
-2022; Dao et al.
-2024), and computer vision (Liu et al.
-2023; Kim et al.
-2024).
-In the graph modeling field, existing RAG efforts have primarily focused on static and text-attributed graphs to enhance the generation capabilities of Large Language Models, supporting graph-related tasks such as code summarization (Liu et al.
-2021) and textual graph question answering (He et al.
-2024; Hu et al.
-2024).
+Related studies have demonstrated the great potential of RAG in various tasks such as language processing (Karpukhin et al., 2020; Jiang et al., 2023), recommendation systems (Ye et al., 2022; Dao et al., 2024), and computer vision (Liu et al., 2023; Kim et al., 2024).
+In the graph modeling field, existing RAG efforts have primarily focused on static (Mao et al., 2023) and text-attributed graphs to enhance the generation capabilities of Large Language Models (LLMs), supporting graph-related tasks such as code summarization (Liu et al., 2021) and textual graph question answering (He et al., 2024; Hu et al., 2024; Jiang et al., 2024).
 However, exploiting RAG techniques for dynamic graphs and graphs without textual information remains largely unexplored.
-3 Preliminaries
+3. Preliminaries
 In this section, we introduce the sequence modeling of dynamic graphs and the problem formulation.
@@
 We denote a dynamic graph as $G=(V,E,F,\mathcal{T})$ comprising a set of nodes $V$, edges $E$, a node feature matrix $F$ if available, and a time domain $\mathcal{T}$.
-To map a dynamic graph into sequences, we follow SimpleDyG (Wu, Fang, and Liao 2024).
+To map a dynamic graph into sequences, we follow SimpleDyG (Wu et al., 2024).
 Specifically, let $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ denote the set of training samples, where each sample is a pair $(x_{i},y_{i})$, representing the input and output sequences for a target node $v_{i}\in V$.
-The input $x_{i}$ is a chronologically ordered sequence of nodes that have historically interacted with $v_{i}$, while the output $y_{i}$ is the ground truth interactions that occurs following the sequence $x_{i}$.
+The input $x_{i}$ is a chronologically ordered sequence of nodes that have historically interacted with $v_{i}$, while the output $y_{i}$ is the ground truth interactions that occur following the sequence $x_{i}$.
 In notations, we have
@@
 The augmented input $\{R_{q},x_{q}\}$ is designed to enhance the predictions of future events in $y_{q}$.
-4 Proposed Model: RAG4DyG
-Our proposed RAG4DyG framework, depicted in Fig.
-2, first trains a generative sequence model for dynamic graphs in Fig.
-2(a).
-We use SimpleDyG (Wu, Fang, and Liao 2024) as the backbone for this task.
-As described in Sect.
-3, a dynamic graph $G$ is represented as sequences of node interaction records, which are then used to train a Transformer-based sequence model.
-Next, given a query sequence $x_{q}$ and a retrieval pool $D$, a time- and context-aware retriever is designed to retrieve demonstrations for $x_{q}$ from $D$, as shown in Fig.
-2(b).
-The retriever fine-tunes its encoder (i.e., the sequence model) to jointly optimize two contrastive losses: a time-aware loss that employs a time decay function, and a context-aware loss that employs sequence augmentation.
-(Sect.
-4.1)
-Finally, given top-K demonstrations retrieved from the retrieval pool, we design a graph fusion module as illustrated in Fig.
-2(c).
-Specifically, we fuse the K demonstrations into a summary graph, leveraging the graph structures inherent to the retrieved demonstrations.
-The summary graph is further encoded by a GNN, serving as an augmented context that is prepended to the original query sequence.
-The augmented sequence is subsequently input to the sequence model to predict future events.
-(Sect.
-4.2)
-4.1 Time- and Context-Aware Retriever
+4. Proposed Model: RAG4DyG
+The RAG4DyG framework enhances dynamic graph modeling by incorporating retrieval-augmented generation techniques to improve predictive accuracy and adaptability.
+As illustrated in Fig.2, it first adopts SimpleDyG (Wu et al., 2024) to model dynamic graphs as sequences of node interactions, leveraging a Transformer-based model to capture temporal dependencies and predict future interactions (Fig.2(a)).
+To enrich the modeling process, a time- and context-aware retriever retrieves relevant demonstrations from a retrieval pool $D$ for a given query sequence $x_{q}$.
+This retriever optimizes two contrastive objectives: a time-aware loss, which employs a temporal decay function $\mu(t_{q},t_{p})$ to prioritize temporally relevant samples, and a context-aware loss, which utilizes sequence augmentation techniques such as masking and cropping to capture structural patterns (Fig.2(b), Sec.4.1).
+Once the top-$K$ demonstrations are retrieved, they are fused into a summary graph $G_{fus}$, which captures the underlying structural relationships among the retrieved samples.
+A GNN then processes this graph to generate an enriched representation that is prepended to the query sequence, providing additional context for improved event prediction (Fig.2(c), Sec.4.2).
+By integrating retrieval and graph fusion, RAG4DyG effectively incorporates temporal and contextual relevance, surpassing existing methods in both transductive and inductive dynamic graph scenarios.
+4.1. Time- and Context-Aware Retriever
 Unlike NLP retrievers, dynamic graph retrieval requires consideration of temporal proximity alongside contextual relevance.
@@
 For each query sequence $x_{q}$, we annotate its positive sample $x_{p}^{+}$ from the pool $D$ based on their contextual similarity.
+We leave the detailed annotation process in Sec.
+5.1.
 Specifically, we adopt the sequence model pre-trained in Fig.
@@
 Given a query sequence $x_{q}$ and a candidate sequence $x_{p}\in D$, we define their contextual similarity as the dot product of their representations:
-$s(x_{q},x_{p})=f(x_{q})^{\top}f(x_{p}),$ (3)
-where $f(\cdot)$ denotes our encoder. We leave the detailed annotation process in Appendix A.
+(3) $s(x_{q},x_{p})=f(x_{q})^{\top}f(x_{p}),$
+where $f(\cdot)$ denotes our encoder.
 Time-aware Contrastive Learning.
@@
 Consequently, we utilize a time decay function to account for temporal proximity between the query and candidate sequences, as follows.
-$\mu(t_{q},t_{p})=\exp(-\lambda|t_{q}-t_{p}|),$ (4)
+(4) $\mu(t_{q},t_{p})=\exp(-\lambda|t_{q}-t_{p}|),$
 where $t_{q}$ and $t_{p}$ represent the last interaction time in the query and candidate sequences [^1] In the annotated training data, the query time $t_{q}$ may precede the candidate time $t_{p}$.
@@
 By using this time decay function, we assign higher importance to the candidates that are temporally closer to the query.
-To effectively capture the temporal dynamics of the graph, we incorporates temporal proximity to reweigh the contextual similarity in the contrastive loss:
-$h(x_{q},x_{p})$ $=s(x_{q},x_{p})\mu(t_{q},t_{p}).$ (5)
+To effectively capture the temporal dynamics of the graph, we incorporate temporal proximity to reweigh the contextual similarity in the contrastive loss:
+(5) $h(x_{q},x_{p})$ $=s(x_{q},x_{p})\mu(t_{q},t_{p}).$
 Subsequently, we adopt in-batch negative sampling based on the following training objective:
-$\mathcal{L}_{\text{tcl}}$ $=-\log\frac{\exp(h(x_{q},x_{p}^{+}))/\tau}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp(h(x_{q},x_{j}))/\tau},$ (6)
+(6) $\mathcal{L}_{\text{tcl}}$ $=-\log\frac{\exp(h(x_{q},x_{p}^{+}))/\tau}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp(h(x_{q},x_{j}))/\tau},$
 where $x_{p}^{+}$ denotes the positive sample of $x_{q}$, $N$ is the batch size, and $\tau$ is the temperature parameter.
@@
 To better capture the inherent contextual pattern, we further adopt context-aware contrastive learning with data augmentations.
-For each sequence, we apply two types of augmentations: masking and cropping, which are widely used for sequence modeling (Devlin et al.
-2019; Xie et al.
-2022).
+For each sequence, we apply two types of augmentations: masking and cropping, which are widely used for sequence modeling (Devlin et al., 2019; Xie et al., 2022).
 The masking operator randomly replaces a portion of the tokens in the sequence with a special masking token.
@@
 Given a sequence $x_{q}$ and its two distinct augmented views $x^{\prime}_{q}$ and $x^{\prime\prime}_{q}$, the contrastive loss is defined as:
-$\mathcal{L}_{\text{ccl}}=-\log\frac{\exp(s(x^{\prime}_{q},x^{\prime\prime}_{q})/\tau)}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp s(x^{\prime}_{q},x^{\prime}_{j})/\tau},$ (7)
+(7) $\mathcal{L}_{\text{ccl}}=-\log\frac{\exp(s(x^{\prime}_{q},x^{\prime\prime}_{q})/\tau)}{\sum_{j=1}^{2N}{\mathds{1}}_{j\neq q}\exp s(x^{\prime}_{q},x^{\prime}_{j})/\tau},$
 where $\tau$ is the temperature, $N$ is the batch size and $\mathds{1}$ is an indicator function.
-Training and Inference for Retrieval.
-The training objective of our retrieval model is defined as:
-$\mathcal{L}_{\text{ret}}=\mathcal{L}_{\text{tcl}}+\alpha\mathcal{L}_{\text{ccl}}$ (8)
-where $\alpha$ is a coefficient that controls the balance between the two losses.
+Training and Inference for Retrieval. The training objective of our retrieval model is defined as:
+(8) $\mathcal{L}_{\text{ret}}=\mathcal{L}_{\text{tcl}}+\alpha\mathcal{L}_{\text{ccl}},$
+where $\alpha$ is a coefficient that balances between the two losses.
 During testing, we utilize the updated sequence model to extract sequence representations and perform demonstration ranking based on the contextual similarity between the query and candidates.
-4.2 Graph Fusion-based Generator
-After retrieval, we obtain the top-K demonstrations $R_{q}=\{(x_{k},y_{k})\}_{k=1}^{K}$ for the query $x_{q}$.
-A straightforward approach is concatenating them with the query sequence.
+4.2. Graph Fusion-based Generator
+After the retrieval process, we obtain the top-K demonstrations $R_{q}=\{(x_{k},y_{k})\}_{k=1}^{K}$ for the query $x_{q}$.
+A straightforward approach is to directly concatenate them with the query sequence and input to a sequence generation model for prediction.
 However, this can lead to a lengthy context that limits the model’s prediction capabilities.
@@
 Thus, we first fuse the demonstrations into a summary graph, process it using a GNN, and then prepend the graph readout from the GNN to the query for subsequent generation.
-Graph Fusion.
 […]
-$e_{\text{fus}}$ $=\mathtt{MeanPooling}(\text{GCN}(G_{\text{fus}})),$ (9)
-$\tilde{x}_{q}$ $=[e_{\text{fus}}\parallel x_{q}],$ (10)
+(9) $e_{\text{fus}}$ $=\mathtt{MeanPooling}(\text{GCN}(G_{\text{fus}})),$
+(10) $\tilde{x}_{q}$ $=[e_{\text{fus}}\parallel x_{q}],$
 where $e_{\text{fus}}$ is the fused graph representation, and $\tilde{x}_{q}$ is the retrieval-augmented sequence.
@@
 Training and Inference.
-We adopt the same sequence model with the same training objective (Wu, Fang, and Liao 2024) as in Fig.
+We adopt the same sequence model with the same training objective (Wu et al., 2024) as in Fig.
 2(a).
 During training, we freeze the parameters of the sequence model, except for the output layer which is updated along with the GCN parameters used for graph fusion.
-During testing, we first apply the retriever model to retrieve top-K demonstrations for each query as introduced in Sect.
+During testing, we first apply the retriever model to retrieve top-K demonstrations for each query as introduced in Sec.
 4.1.
-Then we perform graph fusion on these demonstrations and concatenate the fused graph representation with the query sequence as illustrated in Eqs.
+Then we perform graph fusion on these demonstrations and concatenate the fused graph representation with the query sequence as illustrated in Eq.
 (9) and (10).
-The concatenated sequence is subsequently fed into the trained sequence model for link prediction.
-5 Experiment
+The concatenated sequence is subsequently fed into the trained model for link prediction.
+5. Experiment
 In this section, we empirically evaluate the proposed model RAG4DyG compared to state-of-the-art methods and conduct a detailed analysis of the performance.
-5.1 Experimental Setup
-Datasets.
-We evaluate the performance of the proposed model on three datasets from different domains including the communication network UCI (Panzarasa, Opsahl, and Carley 2009), the citation network Hepth (Leskovec, Kleinberg, and Faloutsos 2005), and the multi-turn task-oriented conversation dataset MMConv (Liao et al.
-2021).
-We follow the same dataset preprocessing as SimpleDyG (Wu, Fang, and Liao 2024).
-It is important to note that the ML-10M dataset was not used because it is not suitable for RAG; we observed no performance enhancement even when the query was augmented with ground-truth demonstrations.
-We summarize the statistics of the three datasets in Table 1.
-Dataset UCI Hepth MMConv
-Domain Social Citation Conversation
-# Nodes 1,781 4,737 7,415
-# Edges 16,743 14,831 91,986
-Table 1: Dataset statistics.
-Method UCI Hepth MMConv
-Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard
-DySAT 0.009±0.003 0.010±0.003 0.010±0.001 0.008±0.004 0.007±0.002 0.005±0.001 0.108±0.089 0.102±0.085 0.095±0.080
-EvolveGCN 0.072±0.046 0.064±0.045 0.032±0.026 0.008±0.002 0.009±0.004 0.007±0.002 0.050±0.015 0.051±0.021 0.032±0.017
-DyRep 0.009±0.008 0.011±0.018 0.010±0.005 0.009±0.006 0.031±0.024 0.010±0.006 0.156±0.054 0.140±0.057 0.067±0.025
-JODIE 0.018±0.019 0.022±0.023 0.012±0.009 0.010±0.008 0.031±0.021 0.011±0.008 0.052±0.039 0.041±0.016 0.032±0.022
-TGAT 0.022±0.004 0.061±0.007 0.020±0.002 0.011±0.007 0.034±0.023 0.011±0.006 0.118±0.004 0.089±0.033 0.058±0.021
-TGN 0.014±0.010 0.041±0.017 0.011±0.003 0.011±0.006 0.030±0.012 0.008±0.001 0.085±0.050 0.096±0.068 0.066±0.038
-TREND 0.083±0.015 0.067±0.010 0.039±0.020 0.010±0.008 0.031±0.003 0.010±0.002 0.134±0.030 0.116±0.020 0.060±0.018
-GraphMixer 0.097±0.019 0.104± 0.013 0.042 ±0.005 0.009±0.002 0.011±0.008 0.010±0.003 0.206±0.001 0.172±0.029 0.085±0.016
-SimpleDyG 0.109±0.014 0.104±0.010 0.092±0.014 0.013±0.006 0.035±0.014 0.013±0.006 0.170±0.010 0.184±0.012 0.169±0.010
-RAG4DyG 0.111±0.013 0.122±0.014 0.097±0.010 0.019±0.002 0.045±0.003 0.019±0.002 0.194±0.005 0.208±0.005 0.194±0.005
-Table 2: Performance comparison for dynamic link prediction. (Best results are bolded; runners-up are underlined.)
-Baselines.
-We compare RAG4DyG with the state-of-the-art approaches of dynamic graph modeling.
-(1) discrete-time approaches: DySAT (Sankar et al.
-2020) and EvolveGCN (Pareja et al.
-2020); (2) continuous-time approaches: DyRep (Trivedi et al.
-2019), JODIE (Kumar, Zhang, and Leskovec 2019), TGAT (Xu et al.
-2020), TGN (Rossi et al.
-2020), TREND (Wen and Fang 2022), GraphMixer (Cong et al.
-2022) and SimpleDyG (Wu, Fang, and Liao 2024).
-The details of the baselines can be found in Appendix B.
-Implementation Details.
-Following the method outlined in (Cong et al.
-2022; Wu, Fang, and Liao 2024), we represent the dynamic graph as an undirected graph.
-We split all datasets into training, validation, and test sets based on temporal sequence same as SimpleDyG (Wu, Fang, and Liao 2024).
+5.1. Experimental Setup
+5.1.1. Datasets.
+We evaluate the performance of the proposed model on six datasets from different domains, including a communication network UCI (Panzarasa et al., 2009), a citation network Hepth (Leskovec et al., 2005), a multi-turn task-oriented conversation dataset MMConv (Liao et al., 2021), a behavioral interaction network Wikipedia (Huang et al., 2024), an email network Enron (Zhang et al., 2024), and a hyperlink network Reddit (Kumar et al., 2018).
+We summarize the statistics of these datasets in Table 1.
+We follow the preprocessing steps of SimpleDyG to map the dynamic graphs into sequences with special tokens (Wu et al., 2024).
+Notably, the Hepth and Reddit datasets exhibit an inductive nature, as they contain previously unseen nodes with no historical interactions.
+The details of the Wikipedia, Enron and Reddit datasets are provided below, while information on the remaining datasets can be found in SimpleDyG (Wu et al., 2024).
+  • Wikipedia (Huang et al., 2024): This dataset captures the co-editing activity on Wikipedia pages over one month.
+It is a bipartite interaction network in which editors and wiki pages serve as nodes.
+Each edge corresponds to an interaction where a user edits a page at a specific timestamp.
+To facilitate temporal sequence alignment, the dataset is divided into 16 time steps based on the timestamps of the interactions.
+  • Enron (Zhang et al., 2024): This dataset represents the email communications among employees of Enron Corporation over three years (1999–2002).
+Nodes represent employees, while edges correspond to emails exchanged between them ordered by the sending timestamps of the emails.
+For temporal sequence alignment, we split the dataset into 17 time steps based on the timestamps.
+  • Reddit (Kumar et al., 2018): This dataset represents a subreddit-to-subreddit hyperlink network, derived from timestamped posts containing hyperlinks between subreddits.
+We focus on hyperlink data within the body of posts, covering the period from 2016 to 2017.
+The dataset is divided into 12 time steps for temporal sequence alignment based on the post timestamps.
+Table 1. Dataset statistics.
+Dataset UCI Hepth MMConv Wikipedia Enron Reddit
+Domain Social Citation Conversation Behavior Social Hyperlink
+# Nodes 1,781 4,737 7,415 9,227 42,711 11,901
+# Edges 16,743 14,831 91,986 157,474 797,907 62,919
+Table 2.
+Performance comparison for dynamic link prediction with mean and standard deviation across 10 runs.
+Best results are bolded; runners-up are underlined.
+* indicates that our model significantly outperforms the best baseline based on the two-tail $t$-test ($p<0.05$).
+Datasets Models DySAT EvolveGCN DyRep JODIE TGAT TGN TREND GraphMixer IDOL SimpleDyG RAG4DyG
+UCI Recall@5 0.009±0.003 0.072±0.046 0.009±0.008 0.018±0.019 0.022±0.004 0.014±0.010 0.083±0.015 0.097±0.019 0.093±0.029 0.109±0.014 0.111±0.013
+NDCG@5 0.010±0.003 0.064±0.045 0.011±0.018 0.022±0.023 0.061±0.007 0.041±0.017 0.067±0.010 0.104±0.013 0.075±0.022 0.104±0.010 0.122*±0.014
+Jaccard 0.010±0.001 0.032±0.026 0.010±0.005 0.012±0.009 0.020±0.002 0.011±0.003 0.039±0.020 0.042±0.005 0.014±0.002 0.092±0.014 0.097±0.010
+Hepth Recall@5 0.008±0.004 0.008±0.002 0.009±0.006 0.010±0.008 0.011±0.007 0.011±0.006 0.010±0.008 0.009±0.002 0.007±0.002 0.013±0.006 0.019*±0.002
+NDCG@5 0.007±0.002 0.009±0.004 0.031±0.024 0.031±0.021 0.034±0.023 0.030±0.012 0.031±0.003 0.011±0.008 0.011±0.003 0.035±0.014 0.045*±0.003
+Jaccard 0.005±0.001 0.007±0.002 0.010±0.006 0.011±0.008 0.011±0.006 0.008±0.001 0.010±0.002 0.010±0.003 0.006±0.001 0.013±0.006 0.019*±0.002
+MMConv Recall@5 0.108±0.089 0.050±0.015 0.156±0.054 0.052±0.039 0.118±0.004 0.085±0.050 0.134±0.030 0.206±0.001 0.169±0.006 0.170±0.010 0.194±0.005
+NDCG@5 0.102±0.085 0.051±0.021 0.140±0.057 0.041±0.016 0.089±0.033 0.096±0.068 0.116±0.020 0.172±0.029 0.115±0.039 0.184±0.012 0.208*±0.005
+Jaccard 0.095±0.080 0.032±0.017 0.067±0.025 0.032±0.022 0.058±0.021 0.066±0.038 0.060±0.018 0.085±0.016 0.015±0.002 0.169±0.010 0.194*±0.005
+Wikipedia Recall@5 0.003±0.005 0.012±0.01 0.003±0.002 0.017±0.005 0.006±0.004 0.016±0.018 0.022±0.012 0.010±0.007 0.022±0.008 0.356±0.016 0.369*±0.006
+NDCG@5 0.002±0.003 0.008±0.007 0.002±0.002 0.015±0.003 0.005±0.005 0.015±0.022 0.016±0.018 0.007±0.006 0.015±0.005 0.398±0.03 0.389±0.008
+Jaccard 0.001±0.001 0.004±0.004 0.001±0.001 0.007±0.002 0.002±0.002 0.007±0.009 0.007±0.021 0.004±0.002 0.004±0.001 0.320±0.027 0.328±0.007
+Enron Recall@5 0.002±0.004 0.004±0.011 0.021±0.001 0.005±0.005 0.020±0.002 0.001±0.001 0.023±0.003 0.021±0.002 0.024±0.014 0.094±0.005 0.100*±0.003
+NDCG@5 0.001±0.002 0.007±0.020 0.036±0.002 0.061±0.039 0.036±0.001 0.003±0.001 0.027±0.001 0.037±0.001 0.025±0.011 0.114±0.005 0.119*±0.004
+Jaccard 0.001±0.001 0.003±0.009 0.019±0.001 0.011±0.007 0.020±0.001 0.001±0.001 0.012±0.001 0.020±0.002 0.008±0.003 0.068±0.003 0.071*±0.002
+Reddit Recall@5 0.001±0.002 0.006±0.002 0.019±0.004 0.013±0.003 0.001±0.001 0.001±0.001 0.002±0.003 0.001±0.001 0.003±0.002 0.101±0.019 0.119*±0.006
+NDCG@5 0.001±0.002 0.012±0.003 0.020±0.004 0.015±0.002 0.001±0.001 0.002±0.001 0.003±0.002 0.003±0.001 0.005±0.003 0.134±0.012 0.143±0.005
+Jaccard 0.001±0.001 0.003±0.001 0.013±0.004 0.007±0.002 0.001±0.001 0.001±0.002 0.001±0.001 0.002±0.001 0.002±0.001 0.088±0.012 0.096±0.003
+5.1.2. Baselines.
+We compare our model RAG4DyG with the state-of-the-art dynamic graph models, which include (1) discrete-time approaches: DySAT (Sankar et al., 2020) and EvolveGCN (Pareja et al., 2020); (2) continuous-time approaches: DyRep (Trivedi et al., 2019), JODIE (Kumar et al., 2019), TGAT (Xu et al., 2020), TGN (Rossi et al., 2020), TREND (Wen and Fang, 2022), GraphMixer (Cong et al., 2022), IDOL (Zhu et al., 2024) and SimpleDyG (Wu et al., 2024).
+  • DySAT (Sankar et al., 2020) utilizes self-attention mechanisms to capture both structural and temporal patterns in dynamic graphs through discrete-time snapshots.
+  • EvolveGCN (Pareja et al., 2020) leverages recurrent neural networks to model the evolution of the parameters of a graph convolutional network over discrete time steps.
+  • DyRep (Trivedi et al., 2019) models dynamic graphs in continuous time by incorporating both temporal point processes and structural dynamics to capture interactions and node dynamics.
+  • JODIE (Kumar et al., 2019) focuses on user and item embedding trajectories over continuous time, predicting future interactions by modeling user and item embeddings jointly.
+  • TGAT (Xu et al., 2020) employs temporal graph attention layers and time encoding to capture temporal dependencies and structural information for dynamic graphs.
+  • TGN (Rossi et al., 2020) combines GNNs with memory modules to maintain node states over continuous time, effectively learning from dynamic interactions.
+  • TREND (Wen and Fang, 2022) integrates temporal dependencies based on the Hawkes process and GNNs to learn the dynamics of graphs.
+  • GraphMixer (Cong et al., 2022) introduces a novel architecture that leverages MLP-mixer to learn link-encoder and node encoder for evolving graphs in continuous time.
+  • IDOL (Zhu et al., 2024) is a contrastive learning-based model tailored for dynamic graph representation learning.
+It utilizes a Personalized PageRank-based algorithm to incrementally update the node embedding and adopt a topology-monitorable sampling method to generate contrastive pairs for efficient training.
+  • SimpleDyG (Wu et al., 2024) reformulates the dynamic graph modeling as a sequence modeling task and mapped the dynamic interactions of target nodes as sequences with specially designed tokens.
+It simplifies dynamic graph modeling without complex architectural changes to effectively capture temporal dynamics.
+5.1.3. Implementation Details.
+Following the method outlined in (Cong et al., 2022; Wu et al., 2024), we represent the dynamic graph as an undirected graph.
+We split all datasets into training, validation, and test sets based on temporal sequence same as SimpleDyG (Wu et al., 2024).
 Given ${T}$ timesteps in each dataset, the data at the final timestep ${T}$ is used as the testing set, the data at ${T-1}$ is served as the validation set, and the remaining data from earlier timesteps is used for training.
 All training data including the retrieval pool for the retriever and generator is drawn from this training data split.
-For retrieval augmented generation model training, we first train SimpleDyG without augmentation using the reported parameter in this paper.
+For retrieval augmented generation model training, we first train SimpleDyG without augmentation using the finetuned parameters.
 Then we fix the parameters of SimipleDyG except for the last linear layer and fine-tune them with the GCN model.
@@
 The number of demonstrations is 7 for all datasets.
-Hyper-parameter tuning is performed on the validation set and we provide the settings in Appendix D.
-Evaluation Metrics.
-We evaluate the performance upon three metrics Recall@5, NDCG@5 and Jaccard (Wu, Fang, and Liao 2024).
-Recall@5 and NDCG@5 are widely used in ranking tasks measuring the quality of top-ranked predictions (Wang et al.
-2019).
-Recall@5 assesses the proportion of relevant nodes in the top 5 predictions, while NDCG@5 evaluates the ranking quality by considering the positions of the relevant nodes.
-Jaccard (Jaccard 1901) measures the similarity between the predicted and ground truth sequences by comparing the intersection and union of the two sets.
-5.2 Performance Comparison
-We evaluate the performance of RAG4DyG for the dynamic link prediction task, and the results compared to the state-of-the-art baselines are presented in Table 2. We make the following observations.
-The proposed RAG4DyG generally outperforms all baselines across different datasets under the three metrics.
+To facilitate retrieval model training, we regard the samples in the training dataset as our retrieval pool $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ where each pair $(x_{i},y_{i})$ represents the historical sequence and its corresponding target sequence.
+Specifically, $x_{i}$ is the input sequence before the last time step and $y_{i}$ is the output sequence at the last time step.
+We annotate demonstrations based on the Jaccard similarity between the output sequences among all the pairs in $D$.
+The Jaccard similarity of two output sequences, $y_{i}$ and $y_{j}$, is given by
+(11) $r(y_{i},y_{j})=\frac{|y_{i}\cap y_{j}|}{|y_{i}\cup y_{j}|}.$
+To control the quality of annotated data, we set a threshold of 0.8 to select highly similar demonstrations for each sample.
+These filtered annotations are then used to train the retriever model.
+The number of training samples for the UCI, Hepth, MMConv, Wikipedia, Enron and Reddit datasets are 9 578, 8 250, 10 762, 162 408, 2 510 666 and 185 764, respectively.
+We ran all the experiments on a Nvidia L40 GPU and tuned the hyper-parameters for all the methods based on the validation set.
+For all the baselines, we tuned the models based on the hyper-parameters reported in their papers.
+For our RAG4DyG method, we set the number of layers, heads and dimensions of hidden states for the backbone SimpleDyG to $(6,8,768)$, $(12,2,256)$, $(2,2,256)$, $(2,6,768)$, $(2,6,768)$, and $(2,8,512)$ across the six datasets.
+The time decay rate $\lambda$ for the retrieval model in Eq.
+(4) was tuned according to the time granularity of different datasets, with days for the UCI dataset, months for the Hepth dataset, turns for the MMConv dataset, and hours for other datasets.
+We explored a range of values $\lambda=\{10^{-4},10^{-3},10^{-2},10^{-1},1,10,100\}$, ultimately selecting $\lambda=10^{-4}$ for UCI, $\lambda=0.1$ for Hepth, $\lambda=10$ for MMConv and Enron, and $\lambda=1$ for Wikipedia and Reddit datasets.
+The coefficient $\alpha$ in the loss function in Eq.
+(8) was tuned across $\{0.2,0.4,0.6,0.8,1\}$, resulting in final values of $\alpha=1$ for UCI and MMConv, $\alpha=0.4$ for Hepth, and $\alpha=0.2$ for Wikipedia, Enron and Reddit.
+Additional parameter settings for the three datasets are as follows: the temperature $\tau$ in the two contrastive learning tasks for all datasets was set to $\tau=0.1$, the batch size of the retriever model for all datasets was set to $N=128$, and the masking and cropping portions in context-aware contrastive learning were set to 0.8, 0.8, 0.8, 0.6, 0.6, 0.2 and 0.4, 0.6, 0.6, 0.8, 0.8, 0.8 across the six datasets, respectively.
+5.1.4. Evaluation Metrics.
+Inspired by SimpleDyG (Wu et al., 2024), we assess the performance of our approach and baselines using three key metrics: Recall@5, NDCG@5, and Jaccard (Wu et al., 2024).
+Recall@5 and NDCG@5 are commonly employed in ranking tasks to evaluate the quality of top-ranked predictions (Wang et al., 2019).
+Specifically, Recall@5 measures the proportion of relevant nodes that appear among the top five predictions, while NDCG@5 considers the ranking positions of the relevant nodes to provide a more nuanced assessment of ranking quality.
+Additionally, the Jaccard index (Jaccard, 1901) quantifies the similarity between the predicted and ground truth sequences by calculating the ratio of their intersection to their union.
+Figure 3. Ablation study for retrieval results.
+Figure 4. Ablation study for link prediction results.
+5.2. Performance Comparison
+We assess the performance of RAG4DyG on the dynamic link prediction task, with the results benchmarked against state-of-the-art baselines, as shown in Table 2.
+Our analysis reveals the following key observations.
+First of all, the proposed RAG4DyG generally outperforms all baselines across different datasets under the three metrics.
 In particular, compared to SimpleDyG, which is also our backbone, RAG4DyG consistently shows superior performance, highlighting the effectiveness of our retrieval-augmented generation framework.
@@
 This indicates that RAG4DyG can better model the temporal and contextual relationships due to the specific design of the retriever and generator.
-Figure 3: Ablation study for retrieval results.
-Second, RAG4DyG exhibits significant advantages in inductive scenarios such as the Hepth dataset.
-This setting is particularly challenging because it involves nodes not appearing during the training phase, requiring the model to generalize to entirely new structures and relationships.
+Generally speaking, the performance of SimpleDyG and RAG4DyG which reformulate the dynamic graph link prediction as a sequence generation task show promising performance compared with node pair ranking-based baselines, especially on the Wikipedia dataset which contains a higher frequency of repeated interaction behaviors.
+This characteristic makes sequence-based models more effective, as they can leverage the temporal consistency and recurrent patterns in the data to better capture the underlying dynamics of the graph.
+Second, RAG4DyG exhibits significant advantages in inductive scenarios such as the Hepth and Reddit datasets.
+This setting is particularly challenging because it involves nodes not seen during training, requiring the model to generalize to entirely new structures and relationships.
 RAG4DyG’s success is attributed to its retrieval-augmented mechanism, which enhances the model’s ability to generalize by providing rich contextual information relevant to the new, unseen nodes.
 Unlike models that rely solely on the immediate neighborhood or predefined structures, RAG4DyG dynamically adapts to the new nodes, ensuring that the predictions are guided by the most relevant and similar historical data.
-5.3 Model Analysis
+5.3. Model Analysis
 We analyze the behavior of our model RAG4DyG in several aspects, including an ablation study, an investigation of the effectiveness of different retrieval methods, and an analysis of parameter sensitivity.
-Ablation Study.
+Table 3. Retrieval performance of various retrieval methods.
+Method UCI Wikipedia Hepth Reddit
+HR@1 HR@3 HR@7 HR@1 HR@3 HR@7 HR@1 HR@3 HR@7 HR@1 HR@3 HR@7
+BM25 0.100 0.136 0.200 0.369 0.405 0.488 - - - - - -
+Jaccard 0.100 0.109 0.146 0.369 0.445 0.430 - - - - - -
+RAG4DyG 0.100 0.155 0.164 0.369 0.455 0.523 0.025 0.049 0.073 0.180 0.218 0.228
+“-” denotes that the method is unable to perform retrieval. The reason is explained in the corresponding description of this table in Sec. 5.3.
+Table 4. Generative performance of various retrieval methods.
+Method UCI Wikipedia Hepth Reddit
+Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard
+BM25 0.111±0.007 0.121±0.009 0.093±0.004 0.368±0.01 0.389±0.012 0.325±0.01 - - - - - -
+Jaccard 0.104±0.009 0.113±0.011 0.088±0.010 0.368±0.013 0.388±0.014 0.321±0.011 - - - - - -
+RAG4DyG 0.111±0.013 0.122±0.014 0.097±0.010 0.369±0.006 0.389±0.008 0.328±0.007 0.019±0.002 0.045±0.003 0.019±0.002 0.119±0.006 0.143±0.005 0.096±0.003
+GroundTruth 0.121±0.010 0.129±0.010 0.107±0.012 0.390±0.008 0.400±0.007 0.340±0.006 0.028±0.004 0.062±0.007 0.028±0.004 0.121±0.008 0.145±0.008 0.099±0.005
+   See the note in Table 3 for the explanation of “-”.
+5.3.1. Ablation Study.
 To evaluate the effectiveness of different modules in the retrieval model, we compare RAG4DyG with two variants w/o CCL and w/o Decay which exclude the context-aware contrastive learning and time decay component in the retrieval model.
@@
 3 and 4, the full model outperforms the two variants, underscoring the benefits of incorporating context-aware contrastive learning and time decay modulation.
-Figure 4: Ablation study for link prediction results.
-Effect of Different Retrieval Methods.
+Notably, the w/o Decay variant exhibits the worst performance across both tasks, emphasizing the critical role of time decay in capturing temporal relevance and accurately modeling the evolving dynamics of the graph.
+5.3.2. Effect of Different Retrieval Methods.
 To further investigate the effectiveness of the retrieval model, we compare our model with two different retrieval methods, namely, BM25 and Jaccard, in Table 3 and 4.
-In Table 4, we only report NDCG@5 and present the remaining metrics in Appendix C.
-BM25 (Fang, Tao, and Zhai 2004) is an extension of the Term Frequency-Inverse Document Frequency (TF-IDF) model, which calculates a relevance score between the query sequence and each candidate sequence in the retrieval pool.
+BM25 (Fang et al., 2004) is an extension of the Term Frequency-Inverse Document Frequency (TF-IDF) model, which calculates a relevance score between the query sequence and each candidate sequence in the retrieval pool.
 The relevance score is derived from the occurrence frequency of the nodes in the query and the retrieval pool.
-Jaccard (Jaccard 1901) measures the similarity between two sets by comparing the size of their intersection to the size of their union.
-Note that in the citation dataset Hepth, all queries in the test set only contain unseen target nodes that never appear in the retrieval pool and have no historical interactions.
+Jaccard (Jaccard, 1901) measures the similarity between two sets by comparing the size of their intersection to the size of their union.
+Note that in the citation dataset Hepth and hyperlink dataset Reddit, the queries in the test set contain unseen target nodes that never appear in the retrieval pool and have no historical interactions.
 As a result, the BM25 and Jaccard scores between the queries and the candidates in the retrieval pool are always zeros.
@@
 For a query sequence containing only the target node, we can still obtain its representation using the sequence model trained for the retrieval model, and further calculate its contextual similarity with the candidate sequences in the retrieval pool.
-Method UCI Hepth MMConv
-HR@1 HR@3 HR@1 HR@3 HR@1 HR@3
-BM25 0.100 0.136 - - 0.025 0.233
-Jaccard 0.100 0.109 - - 0.007 0.213
-RAG4DyG 0.096 0.153 0.025 0.049 0.034 0.168
-“-” denotes that the method is unable to perform retrieval. The reason is explained in the corresponding description of this table in Sect. 5.3.
-Table 3: Retrieval performance of various retrieval methods.
-Method UCI Hepth MMConv
-BM25 0.121±0.009 - 0.207±0.003
-Jaccard 0.$1$ 0.207±0.005
-RAG4DyG 0.122±0.014 0.045±0.003 0.208±0.005
-GroundTruth 0.129±0.010 0.062±0.007 0.224±0.008
-See the note in Table 3 for the explanation of “-”.
-Table 4: Generative accuracy of various retrieval methods.
-In Table 3, we analyze the retrieval performance of different retrieval methods.
-Our retrieval model shows better performance than other retrieval strategies.
-Notably, in inductive scenarios like the Hepth dataset, BM25 and Jaccard fail to work with new query nodes lacking historical interactions.
+Figure 5. Effect of the number of demonstrations K.
+In Table 3, we analyze the retrieval performance of different methods.
+For transductive scenarios, our retrieval model shows comparable performance to other retrieval strategies.
+Notably, in inductive scenarios like the Hepth and Reddit datasets, BM25 and Jaccard fail to work with new query nodes lacking historical interactions.
 In contrast, our model can handle them effectively and achieve solid performance.
-The high $HR@3$ performance of BM25 and Jaccard on the MMConv dataset can be attributed to the nature of the dialogue dataset, where temporal order is less critical, and certain nodes associated with specific slot values are more discriminative.
 Table 4 shows the generative performance of different retrieval methods in the dynamic link prediction task.
@@
 Our method performs better compared to other retrieval strategies, indicating the effectiveness of contrastive learning in the retrieval model.
-Effect of the Number of Demonstrations K.
+5.3.3. Effect of the Number of Demonstrations K
 To investigate the influence of the number of demonstrations, we conduct experiments across varying values $K\in\{1,3,5,7,9\}$.
@@
 However, including too many cases may introduce more noise, which can harm the performance.
-Figure 5: Effect of the number of demonstrations K.
-Effect of Different Fusion Strategies.
-To further investigate the effectiveness of the fusion strategy for the top-K demonstrations, we conduct experiments with different fusion strategies in Table 5 for K $=7$.
-We report NDCG@5 in Table 4 and present the other metrics in Appendix C.
+Ultimately, we select $K=7$ for all datasets.
+Table 5. Effect of different fusion strategies.
+Fusion strategy UCI Wikipedia Hepth Reddit
+Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard Recall@5 NDCG@5 Jaccard
+Concatenation 0.033±0.019 0.036±0.018 0.029±0.016 0.210±0.019 0.232±0.021 0.206±0.019 0.001±0.002 0.007±0.002 0.002±0.002 0.001±0.001 0.003±0.003 0.001±0.001
+MLP 0.102±0.018 0.106±0.017 0.089±0.016 0.356±0.006 0.371±0.009 0.321±0.007 0.006±0.002 0.015±0.002 0.006±0.002 0.108±0.006 0.132±0.005 0.090±0.003
+GraphFusion 0.111±0.013 0.122±0.014 0.097±0.010 0.369±0.006 0.389±0.008 0.328±0.007 0.019±0.002 0.045±0.003 0.019±0.002 0.119±0.006 0.143±0.005 0.096±0.003
+5.3.4. Effect of Different Fusion Strategies.
+To further investigate the effectiveness of the fusion strategy for the top-K demonstrations, we conduct experiments with different fusion strategies under K $=7$.
 “Concatenation” denotes we directly concatenate the sequences of retrieved demonstrations and prepend them with the query sample sequence and then feed it into the pre-trained SimpleDyG model.
@@
 The superior performance of the “GraphFusion” strategy highlights the importance of considering both the content and the structure of the demonstrations in the fusion process.
-Fusion strategy UCI Hepth MMConv
-Concatenation 0.036±0.018 0.007±0.002 0.110±0.030
-MLP 0.106±0.017 0.015±0.002 0.167±0.006
-GraphFusion 0.122±0.014 0.045±0.003 0.208±0.005
-Table 5: Effect of different fusion strategies.
-6 Conclusion
-In this work, we introduced a novel retrieval augmented framework RAG4DyG for dynamic graph modeling, addressing the limitations of existing approaches that often rely on narrow historical contexts.
-We leverage the RAG technique to broaden the context by incorporating relevant auxiliary information by harvesting high-quality demonstrations for dynamic graph samples and effectively utilizing this demonstration to enhance dynamic graph modeling.
-Our proposed solution involves a time-aware contrastive learning model to identify and retrieve temporally pertinent cases for each query sequence, coupled with a graph fusion strategy to integrate the inherent historical context with extended temporal contexts.
-Extensive experiments on real-world datasets across different domains demonstrated the superior performance of RAG4DyG in dynamic graph modeling.
-Appendix
-Appendix A Retrieval Data Annotation
-To facilitate retrieval model training, we regard the samples in the training dataset as our retrieval pool $D=\{(x_{i},y_{i})\}_{i=1}^{M}$ where each pair $(x_{i},y_{i})$ represents the historical sequence and its corresponding target sequence.
-Specifically, $x_{i}$ is the input sequence before the last time step and $y_{i}$ is the output sequence at the last time step.
-We annotate demonstrations based on the Jaccard similarity of $y_{i}$ among all the pairs in $D$.
-$r(y_{i},y_{j})=\frac{|y_{i}\cap y_{j}|}{|y_{i}\cup y_{j}|}$ (11)
-To control the quality of annotated data, we set a threshold of 0.8 to select highly similar demonstrations for each sample.
-These filtered annotations are then used to train the retriever model.
-The number of training samples for UCI, Hepth and MMConv datasets are 9,578, 8,250 and 10,762, respectively.
-Appendix B More Details about Baselines
-The details of the baselines are as follows:
-  • DySAT (Sankar et al. 2020) utilizes self-attention mechanisms to capture both structural and temporal patterns in dynamic graphs through discrete-time snapshots.
-  • EvolveGCN (Pareja et al. 2020) leverages recurrent neural networks to evolve the graph convolutional network parameters over discrete time steps.
-  • DyRep (Trivedi et al. 2019) models dynamic graphs in continuous time by incorporating both temporal point processes and structural dynamics to capture interactions and node dynamics.
-  • JODIE (Kumar, Zhang, and Leskovec 2019) focuses on user and item embedding trajectories over continuous time, predicting future interactions by modeling user and item embeddings jointly.
-  • TGAT (Xu et al. 2020) employs temporal graph attention layers and time encoding to capture temporal dependencies and structural information for dynamic graphs.
-  • TGN (Rossi et al. 2020) combines GNNs with memory modules to maintain node states over continuous time, effectively learning from dynamic interactions.
-  • TREND (Wen and Fang 2022) integrates temporal dependencies based on Hawkes process and GNN to learn the dynamics of graphs.
-  • GraphMixer (Cong et al. 2022) introduces a novel architecture that leverages MLP-mixer to learn link-encoder and node encoder for evolving graphs in continuous time.
-  • SimpleDyG (Wu, Fang, and Liao 2024) reformulated the dynamic graph modeling as a sequence modeling task and mapped the dynamic interactions of target nodes as sequences with specially designed tokens.
-It simplifies dynamic graph modeling without complex architectural changes to effectively capture temporal dynamics.
-Appendix C More Experimental Results
-For the effect of different retrieval methods for generative performance in Sec.
-5.3, we report the NDCG@5 score in Table 4.
-Here we report the remaining metrics in Table 6.
-The results show that RAG4DyG performs better than BM25 and Jaccard on Recall@5 and Jaccard metrics, which are consistent with the performance of NDCG@5.
-Method UCI Hepth MMConv
-Recall@5 Jaccard Recall@5 Jaccard Recall@5 Jaccard
-BM25 0.111±0.007 0.093±0.004 - - 0.193±0.003 0.192±0.003
-Jaccard 0.104±0.009 0.088±0.010 - - 0.193±0.004 0.192±0.004
-RAG4DyG 0.111±0.013 0.097±0.010 0.019±0.002 0.019±0.002 0.194±0.005 0.194 ±0.005
-GroundTruth 0.121±0.010 0.107±0.012 0.028±0.004 0.028±0.004 0.210±0.008 0.208±0.008
-Table 6: Generative performance of retrieval methods.
-For the effect of different fusion strategies in Sec.
-5.3, the performance of the remaining metrics is shown in Table 7.
-Similarly, the “GraphFusion” strategy outperforms direct concatenation and “MLP”, indicating the effectiveness of the graph structure among demonstrations.
-Fusion strategy UCI Hepth MMConv
-Recall@5 Jaccard Recall@5 Jaccard Recall@5 Jaccard
-Concatenation 0.033±0.019 0.029±0.016 0.001±0.002 0.002±0.002 0.103±0.031 0.103±0.031
-MLP 0.102±0.018 0.089±0.016 0.006±0.002 0.006±0.002 0.156±0.006 0.153±0.006
-GraphFusion 0.111±0.013 0.097±0.010 0.019±0.002 0.019±0.002 0.194±0.005 0.194 ±0.005
-Table 7: Effect of different fusion strategies.
-Appendix D Hyper-parameter Settings
-We run all the experiments on an NVIDIA GPU L40 with the same data splitting and follow the hyper-parameter settings reported in SimpleDyG (Wu, Fang, and Liao 2024) for all the baselines.
-For our RAG4DyG method, we report the hyper-parameters as follows.
-The time decay rate $\lambda$ for the retrieval model in Eq.
-(4) was fine-tuned according to the time granularity of different datasets, with days for the UCI dataset, months for the Hepth dataset, and turns for the MMConv dataset.
-We explored a range of values $\lambda=\{1e-4,1e-3,1e-2,1e-1,1,10,100\}$, ultimately selecting $\lambda=1e-4$ for UCI, $\lambda=0.1$ for Hepth, and $\lambda=10$ for MMConv.
-The coefficient $\alpha$ in the loss function (Eq.
-(8) was tuned across $\{0.2,0.4,0.6,0.8,1\}$, resulting in final values of $\alpha=1$ for UCI and MMConv dataset, $\alpha=0.4$ for Hepth.
-Additional parameter settings for the three datasets are as follows: the temperature $\tau$ in the two contrastive learning tasks is set to $\tau=\{0.1,0.1,0.1\}$; the batch size of the retriever model is $N=\{64,128,128\}$; and the masking and cropping portions in context-aware contrastive learning are $\{0.8,0.8,0.8\}$ and $\{0.4,0.6,0.6\}$, respectively.
+5.3.5. Time Complexity Analysis
+The time complexity of our model RAG4DyG aligns with that of the vanilla Transformer, scaling as $O(n^{2})$, where $n$ denotes the sequence length.
+We measured the training time per epoch across various approaches using the UCI dataset to assess its efficiency.
+The results shown in Table 6 indicate that our model achieves faster or comparable training cost compared to the baseline methods.
+In contrast, approaches that integrate temporal components (e.g., RNNs or self-attention mechanisms) with structural elements (e.g., GNNs or GATs) face significant computational challenges due to the complexity of combining these modules.
+Figure 6. Time efficiency and Recall@5 of different methods.
+6. Conclusion
+In this work, we proposed RAG4DyG, a novel retrieval-augmented framework for dynamic graph modeling that overcomes the limitations of existing approaches by integrating broader temporal and contextual information.
+By leveraging the retrieval-augmented generation paradigm, RAG4DyG retrieves high-quality demonstrations from a retrieval pool and incorporates them effectively into the modeling process.
+The framework includes a time-aware contrastive learning module to prioritize temporally relevant samples and a graph fusion strategy to seamlessly integrate these retrieved demonstrations with the query sequence, enriching the historical context with extended temporal insights.
+Extensive experiments on diverse real-world datasets demonstrate the effectiveness of RAG4DyG in achieving state-of-the-art performance for dynamic graph modeling in both transductive and inductive scenarios.
+Acknowledgements.
+This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041 and T2EP20123-0052).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
+The authors would also like to thank HOANG Thi Linh for her assistance with this work.